How many gallons of a 70% antifreeze solution must be mixed with 60 gallons of a 15% antifreeze solution to make a 60% antifreeze solution?
Answer by mananth(16946)

--------percent ---------------- quantity
Antifreeze I 15 ---------------- 60 gal
Antifreeze II 70 ---------------- x gal
Mixture 60 ---------------- 60 + x gal

15*60+70x=60( 60 + x )
900 + 70 x = 3600 + 60 x
70 x + -60 x = 3600 - -900
10 x = 2700
/ 10
x = 270
270 gal of 70 % Antifreeze II